iYuko (iyuko):

Obtaining the Equation of a Line Lesson 1 on Graphing Come Learn with me!

iYuko (iyuko):

If you know the slope of a line and the y-intercept, we can write an equation of a line in slope-intercept form. Sometimes we are given the slope and a point on the line. We use the information to find the y-intercept. Then we can write an equation of the line.

iYuko (iyuko):

Find the equation of a line with a slope of 4 that passes through the point (3,1). We Are given the fallowing values: M=4, X=3, Y=1. Y=mx+b Y=4x+b We are given that the slope of the line is 4. 1=4(3)+b Since (3,1) is a point on the line, it satifies the equation Subisitute x = 3 and y = 1 into the equation. Solve for b. 1=12+b -11=b Thus the y-intercept is -11. We can now write the equation of the line y=4x-11

iYuko (iyuko):

To find the equation of a line given a point and a slope 1. Sub the given values of x,y, and m into the equation y=mx+b 2. Solve for b 3. Use the values of b and m to write the equation in the form of y=mx+b

OpenStudy (nincompoop):

are you sure?

Nnesha (nnesha):

-2/5 is slope not -2/3

OpenStudy (nincompoop):

given a slope and a point, use the point-slope form \(\large y-y_1 = m (x-x_1) \) then solve for y

Nnesha (nnesha):

nope what about (-2/3)(-3) ?

OpenStudy (nincompoop):

are you paying attention or what?

OpenStudy (nincompoop):

read the information I provided

Nnesha (nnesha):

better to change slope which is in the question -2/5 to -2/3

OpenStudy (nincompoop):

alright I already told you what to use but you keep insisting

Nnesha (nnesha):

what did i say ? that's an example you already solved that if you want -2/5 then you to solve it again (-2/5)(-3) = ?

Nnesha (nnesha):

@nincompoop can't we use slope intercept form ??

iYuko (iyuko):

Find the equation of the line with the slope of \[\rm-\frac{ 2 }{ 3 }\] And passes throught the point (-3,6) We write the values out \[\rm \color{red}{m}=-\frac{ 2 }{ 3 }, x=-3, y=6\]\[\huge\rm y=\color{red}{m}x+\color{blue}{b}\] \[\large\rm 6 = (-\frac{ 2 }{ 3 })(-3) +\color{blue}{ b}\]\[\rm 6=2+b\]\[\rm4=\color{blue}{b}\] The equation of the line is \[\large\rm y=\color{red}{-\frac{ 2 }{ 3 }}x+\color{blue}{4}.\]